# Drop rebound at low Weber number

### v1

@@ Appendix A Defining contact
Figure 9: ($a$) Magnified view near a drop’s contact point during impact. The hollow circle indicates the contact point with the substrate, while the filled circle marks the contact point along a hypothetical substrate at a height corresponding to the experimental spatial resolution $z=0.02R$. This threshold increases the contact time and expands the contact radius $r_{c}$ by $b$. ($b$) Dimensionless contact time $\bar{t}_{c}=t_{c}/t_{\sigma}$ and ($c$) dimensionless maximum contact radius $\beta_{c}=R_{c}/R$ against Weber number for the KM model evaluated at the hypothetical substrate $0.02R$ (gray) and true substrate $0.00R$ (black) for $O\kern-0.45003pth\,=0.03$. Red curves represent predictions from the model of Chevy et al. (2012). Solid lines show predictions for $B\kern-0.45003pto\,=0.019$, and dashed lines show results obtained for $B\kern-0.45003pto\,=0$.
@@ B.2 Finite differences in time
[…]
@@
$-\delta_{t}^{k}\,B_{1}^{k+1}+v^{k+1}=v^{k}-\delta_{t}^{k}\,B\kern-0.50003pto\,;$ (50)
@@
[…]
@@ B.2.1 Angular discretization
The spectral mesh is determined by the maximum order of harmonic modes $L$ in our truncated expansion, using Legendre polynomials of degree 00 to $L$. The angular mesh is naturally defined as the union of ${\theta=0}$ (the south pole) and the zeros of $P_{L}(\cos(\theta))$, which is canonical for the Legendre polynomial basis (Boyd, 2000). A schematic of this angular mesh is shown in figure 3($c$).
@@
[…]
$-\delta_{t}^{k}\,v^{k+1,q}+h^{k+1,q}=h^{k},$ (61)
@@
$-\delta_{t}^{k}\,B_{1}^{k+1,q}+v^{k+1,q}=v^{k}-\delta_{t}^{k}\,B\kern-0.50003pto\,,$ (62)
@@ B.4 The system matrices
Here, $I_{m}$ is the identity matrix of size $m$, and from equation (60) we have
@@
where the first two zero columns represent modes 00 and $1$, absent in equation (60).
@@ C.2 Finding local minima of the error
Assuming the contact area boundary moves continuously, we can reduce the computational cost of solving the discrete optimization problem in equation (66), which would otherwise require evaluating all possible $q$ from 00 to $L$. The error in the angle between the drop and substrate (the quantity to minimize) exhibits clear monotonic behavior around the minimum. Figure 11($a$) shows the normalized error $\bar{E}=2\,\textrm{tan}^{-1}(E)/\pi$ plotted against $q$ for a case where the solution indicates that $q=10$ contact points is the optimal solution, i.e., $m^{k+1}=10$. Our normalization of the error allows us to display both finite and infinite errors in the same plot, with $\bar{E}=1$ for $E=\infty$.

### v2

@@ Appendix A Defining contact
Figure 9: ($a$) Magnified view near a drop’s contact point during impact. The hollow circle indicates the contact point with the substrate, while the filled circle marks the contact point along a hypothetical substrate at a height corresponding to the experimental spatial resolution $z=0.02R$. This threshold increases the contact time and expands the contact radius $r_{c}$ by $b$. ($b$) Dimensionless contact time $\bar{t}_{c}=t_{c}/t_{\sigma}$ and ($c$) dimensionless maximum contact radius $\beta_{c}=R_{c}/R$ against Weber number for the KM model evaluated at the hypothetical substrate $0.02R$ (gray) and true substrate $0.00R$ (black) for $O\kern-0.46252pth\,=0.03$. Red curves represent predictions from the model of Chevy et al. (2012). Solid lines show predictions for $B\kern-0.46252pto\,=0.019$, and dashed lines show results obtained for $B\kern-0.46252pto\,=0$.
@@ B.2 Finite differences in time
$\delta^{k}_{t}\,l\,(l+2)(l-1)\,A^{k+1}_{l}+\left[1+\delta^{k}_{t}\,2(2l+1)(l-1)\,O\kern-0.50003pth\,\right]\,{U}_{l}^{k+1}+\delta^{k}_{t}\,l\,B_{l}^{k+1}=U^{k}_{l},$ (48)
$\text{ for }l=2,\ldots L,$
@@
$\delta_{t}^{k}\,B_{1}^{k+1}+v^{k+1}=v^{k}-\delta_{t}^{k}\,B\kern-0.50003pto\,;$ (50)
@@
$\sum\limits_{l=0}^{L}{B_{l}^{k+1}P_{l}(\cos(\theta))}=0,\qquad\hskip 62.59605pt\text{ for }\theta>\theta_{c};$ (53)
@@ B.2.1 Angular discretization
The spectral mesh is determined by the maximum order of harmonic modes $L$ in our truncated expansion, using Legendre polynomials of degree 0 to $L$. The angular mesh is naturally defined as the union of $\theta=0$} (the south pole) and the zeros of $P_{L}(\cos(\theta))$, which is canonical for the Legendre polynomial basis (Boyd, 2000). A schematic of this angular mesh is shown in figure 3($c$).
@@
$\delta^{k}_{t}\,l\,(l+2)(l-1)\,A^{k+1,q}_{l}+\left[1+\delta^{k}_{t}\,2(2l+1)(l-1)\,O\kern-0.50003pth\,\,\right]\,{U}_{l}^{k+1,q}+\delta^{k}_{t}\,l\,B_{l}^{k+1,q}=U^{k}_{l},$ (60)
$\text{ for }l=2,\ldots L,$
$\delta_{t}^{k}\,v^{k+1,q}+h^{k+1,q}=h^{k},$ (61)
@@
$\delta_{t}^{k}\,B_{1}^{k+1,q}+v^{k+1,q}=v^{k}-\delta_{t}^{k}\,B\kern-0.50003pto\,$, (62)
@@ B.4 The system matrices
Here, $I_{m}$ is the identity matrix of size mm, and from equation (60) we have
@@
where the first two zero columns represent modes 0 and $1$, absent in equation (60).
@@ C.2 Finding local minima of the error
Assuming the contact area boundary moves continuously, we can reduce the computational cost of solving the discrete optimization problem in equation (66), which would otherwise require evaluating all possible $q$ from 0 to $L$. The error in the angle between the drop and substrate (the quantity to minimize) exhibits clear monotonic behavior around the minimum. Figure 11($a$) shows the normalized error $\bar{E}=2\,\textrm{tan}^{-1}(E)/\pi$ plotted against $q$ for a case where the solution indicates that $q=10$ contact points is the optimal solution, i.e., $m^{k+1}=10$. Our normalization of the error allows us to display both finite and infinite errors in the same plot, with $\bar{E}=1$ for $E=\infty$.